OpenStudy (anonymous):

If anyone can do this step-by-step for me I would appreciate it! Please and Thank You! ln(x-5)-ln(x+4)=ln(x-1)-ln(x+2)

OpenStudy (anonymous):

\[\ln(x-5)-\ln(x+4)=\ln(x-1)-\ln(x+2)\implies \frac{x-5}{x+4}=\frac{x-1}{x+2}\]\[\implies x^2-3x-10=x^2+3x-4\implies 6x=-6\implies x=-1\]But x=-1 is not in the domain of the problem, so there is no solution. Answers like this are always disappointing, but I've checked a couple of times, and am pretty sure that's the case.

OpenStudy (anonymous):

Thank You So Much!